SIR, I WANT TO A SOLUTION OF MY QUESTION THAT IS ARE AS UNDER:
42 APPLE IN NUMBER,BUT I WANT TO DIVIDE TEASE APPLE INTO SEVEN PERSONS,BUT CONDITION IS HERE, EVERY PERSONS HAS A OWN SHARE OF APPLE IN FORM OF ODD NUMBER.
THANKS REGARDS
DINESH JANGRA
PANIPAT

Answer by Theo(13342)

i'm not sure you can do it.
suppose only 3 people and 18 apples.
this means 6 apples per person.
can you give each person an odd number of apples.
suppose you gave 1 and 1. the balance is 16 which is an even number.
suppose you gave 3 and 3. same deal.
suppose you gave 5 and 5. the balance is an even number (8)
suppose you gave 7 and 7. the balance is an even number 4
no matter how you slice up an odd number, the balance will always be even.
1 + 3 = 4 from 18 = 14
1 + 5 = 6 from 18 = 12
3 + 5 = 8 from 18 = 10
you cannot get 3 odd numbers to add up to 18.
similarly, i do not believe you can get 7 odd numbers to add up to 42.
every 2 of those number will add up to an even number.
the sum of 3 pairs of even numbers will be an even number.
when you subtract that sum from 42, you will get an even number as your answer.
the 7th person has to be even if all the other 6 persons received an odd number.
this is because an odd number plus an odd number will always be an even number.
that's what i think anyway.
if you find an answer that says that you can do it, by all means, let me know.